Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 02, 1997, pp. 20–28. Submitted: September 10, 1997. Published: December 30, 1997.

COINCIDENCE POINTS OF COMPATIBLE MULTIVALUED MAPPINGS

R.A.RASHWAN AND M.A.AHMED

Abstract. In this paper we establish a coincidence theorem for a generalized contractive type multivalued mappings in complete metric spaces generalizing some results of Constantin [1]. Two examples are given to support our generalization.

1991 AMS Mathematics subject classification Code. 54H25..

Key words and phrases. Coincidence points, compatible mappings, complete metric spaces.

Typeset by AMS-TEX

Department of Mathematics, Faculty of Science, Assiut University, Assiut, EGYPT

Department of mathematics, Faculty of science, Assiut University, Assiut. EGYPT

\begin{matrix}c \\ ◯ \end{matrix}1997 Cameron University